Show all work please!!!

8) Use 5 terms of the binomial formula to approximate the square root of 15.
Answer #1


We look at Binomial expansion of (1 - 2x)^\frac{1}{2} to approximate \sqrt{15} .

We know that

(1 + x)^n = 1 + \frac{n}{1}x + \frac{n(n-1)}{1*2}x^2 + ...

Thus, the expansion of (1 - 2x)^\frac{1}{2} is

                                                      = 1 - x -\frac{1}{2}x^2 - \frac{1}{2}x^3 -\frac{5}{8}x^4 ...

For x=-7 ,

(1 - 2x)^\frac{1}{2}=(1 - 2(-7))^\frac{1}{2}=\sqrt{15}

So

\begin{align*} \sqrt{15}& \sim 1 - (-7) -\frac{1}{2}(7)^2 - \frac{1}{2}(-7)^3 -\frac{5}{8}(-7)^4\\ & =1 +7 -\frac{1}{2}7^2 +\frac{1}{2}7^3 -\frac{5}{8}7^4\\ &=8-24.5+171.5-1500.625\\&=-1345.625 \end{align}